OpenStudy (rea201):

Find the pattern in the set of data. Then write and equation. x values:1,2,3,4,5,6,7 y values:1,3,7,15,31,63,127

OpenStudy (rea201):

@Hero

OpenStudy (rea201):

I know they add ,4,8,16,32,64

OpenStudy (rea201):

@zepdrix

Directrix (directrix):

@Rea201 Double the previous value of y and add what number?

OpenStudy (rea201):

Ohh ok but what would the equation be? I see your multiplying the previous y by 2 and adding 1.

Directrix (directrix):

2^1 - 1 = 2-1 = 1 2^2 - 1 = 4-1 = 3 2^3 - 1 = 8-1 = 7 2^4 - 1 =16-1=15 How would you write 31 in this way? @Rea201

OpenStudy (rea201):

2^5-1